Range of a cannon ball. Suppose one of Napoleon’s cannons had a muzzle speed, vi, of 60 m/s. At what angle should it have been aimed (ignore air resistance) to strike a target 320 m away?


1
Expert's answer
2020-11-10T07:02:54-0500

The length of flying for the body launched at the angle "\alpha" with the initial speed "v_0 = 60m\/s" is (see https://opentextbc.ca/universityphysicsv1openstax/chapter/4-3-projectile-motion/):


"L = \dfrac{v_0^2\sin2\alpha}{g}"

where "g = 9.81m\/s^2" is the gravitational acceleration. Expressing "\alpha", obtain:


"\alpha =\dfrac{1}{2}\arcsin \left( \dfrac{Lg}{v_0^2}\right)"

If "L = 320m", then:


"\alpha =\dfrac{1}{2}\arcsin \left( \dfrac{320\cdot 9.81}{60^2}\right) \approx 30.3\degree"

Answer. "30.3\degree".